# Conformal Prediction with Upper and Lower Bound Models

### v1

@@ 3.1 Exploiting valid bounds in NCP
Recall that $\hat{B}^{l},\hat{B}^{u}$ provide valid lower and upper bounds on the target variable, i.e., $\hat{B}^{l}(X)\,{\leq}\,Y\,{\leq}\,\hat{B}^{u}(X)$ always holds. This suggests a simple strengthening procedure, whereby a prediction interval $\hat{C}$ is strengthened into
@@
As shown by Proposition 3.1, this is without loss of coverage.
@@
$\mathbb{P}\big{(}Y_{N+1}\,{\in}\,\tilde{C}(X_{N+1})\big{)}=\mathbb{P}\big{(}Y_{N+1}\,{\in}\,\hat{C}(X_{N+1})\big{)}=1{-}\alpha.$
{appendixproof}
It suffices to prove that $\mathbb{P}\left(Y\in\tilde{C}(X)\right)=\mathbb{P}\left(Y\in\hat{C}(X)\right)$. First note that
@@
which concludes the proof.
It is important to note that Proposition 3.1 holds irrespective of how the original prediction interval $\hat{C}$ is obtained. Theorem 3.1 shows that, if the calibration step is performed using the NCP framework, then there is no loss of performance whether the strengthening is performed before or after the calibration step.
{theoremrep}
Consider a family of nested prediction sets $\{\hat{C}_{t}\}_{t\in\mathcal{T}}$, where $\mathcal{T}\,{\subseteq}\,\mathbb{R}$, and let $\hat{\tau}$ be obtained following the NCP calibration step as per Eq. (12). Next, define the family of nested strengthened intervals $\{\tilde{C}_{t}\}_{t\in\mathcal{T}}$, where
$\forall x\in\mathbb{R}^{d},\forall t\in\mathcal{T},\tilde{C}_{t}(x)=\hat{C}_{t}(x)\cap\big{[}\hat{B}^{l}(x),\hat{B}^{u}(x)\big{]},$
and let $\tilde{\tau}$ be obtained from Eq. (12). Then, $\hat{\tau}=\tilde{\tau}$ and
$\forall x\in\mathbb{R}^{d},\hat{C}_{\hat{\tau}}(x)\cap[\hat{B}^{l}(x),\hat{B}^{u}(x)]$ $=\tilde{C}_{\tilde{\tau}}(x).$ (16)
@@
which then yields $\hat{\tau}=\tilde{\tau}$ by unicity of the infimum, and $\hat{C}_{\hat{\tau}}(X)\cap[\hat{B}^{l}(X),\hat{B}^{u}(X)]=\tilde{C}_{\tilde{\tau}}(X)$ follows immediately.
Theorem 3.1 allows to decouple the calibration step from the strengthening. Thereby, one can perform the calibration step without access to lower/upper bound predictors, without any loss of coverage. This is particularly valuable when lower/upper bounds are not available during training/calibration due to, e.g., privacy concerns.
@@ 3.2 The CPUL approach
Note that $L^{l}$, $L^{u}$, $U^{l}$, $U^{u}$ do not provide valid lower nor upper bounds on the target variable ($Y$). The motivation for adjusting the initial predictors $\hat{B}^{l},\hat{B}^{u}$ using quantiles $\hat{Q}^{l},\hat{Q}^{u}$ is to account for the heteroskedascitity of $\hat{r}^{l},\hat{r}^{u}$. Also note that the construction in Eq. (19) only requires evaluating the residuals $\hat{r}^{l},\hat{r}^{u}$ on the training set, and extracting their quantiles. This is more efficient than the approach proposed in sesia2023conformal, wherein additional quantile regression models are trained to predict $\hat{r}^{l},\hat{r}^{u}$.
@@
This construction is illustrated in Figure 1. Each family is then conformalized using the NCP procedure on the calibration set, and the model with smallest width is selected.
Figure 1: Illustration of the construction of prediction intervals in CPUL: (a) $\hat{C}^{\text{ll}}_{0}$, (b) $\hat{C}^{\text{uu}}_{0}$, (c) $\hat{C}^{\text{lu}}_{0}$, (d) $\hat{C}^{\text{ul}}_{0}$. Each family of prediction intervals is conformalized using the calibration set, and following the nested CP framework (see Eq. (12)).
@@
Algorithm 1 summarizes the proposed CPUL method, which proceeds as follows. First, the dataset $\mathcal{D}$ is split into training and calibration sets (line 1). Then, lower and upper bound predictors (i.e., $\hat{B}^{l},\hat{B}^{u}$) are trained using the training set (line 2), after which quantiles of the corresponding residuals are evaluated on the training set (line 3). Next, four families of nested intervals are constructed following Eq. (20), each of which is calibrated using the NCP framework (Eq. (12)) and the calibration set $\mathcal{D}_{\text{cal}}$ (line 4). The final step of the algorithm (line 5) selects, among the four models $\hat{C}^{\text{ll}},\hat{C}^{\text{lu}},\hat{C}^{\text{ul}},\hat{C}^{\text{uu}}$, the one with smallest average width over the calibration set (as an estimator of $\mathbb{E}\big{|}\hat{C}(X_{N+1})\big{|}$).
It is important to note that the model selection step is not performed on a per-sample basis. Rather, Algorithm 1 selects one variant among $\hat{C}^{\text{ll}},\hat{C}^{\text{lu}},\hat{C}^{\text{ul}},\hat{C}^{\text{uu}}$, which is then used across the entire test set. For instance, if $\hat{C}^{\text{lu}}_{\tau^{\text{lu}}}$ is selected, then $\hat{C}^{**}\,{=}\,\hat{C}^{\text{lu}}_{\tau^{\text{lu}}}$ and the CPUL prediction interval is $\hat{C}^{**}(X_{N+1})\,{=}\,\hat{C}^{\text{lu}}_{\tau^{\text{lu}}}(X_{N+1})$. Theorem 1 provides theoretical guarantees on the coverage of $\hat{C}^{**}$.
Theorem 1.
Assume that $\{(X_{1},Y_{1}),...,(X_{N+1},Y_{N+1})\}$ are i.i.d. samples, and let $\hat{C}^{**}$ be the CPUL model selected by Algorithm 1. Then
$\mathbb{P}\left(Y_{N+1}\in\hat{C}^{**}(X_{N+1})\right)\geq\frac{1{+}N_{\text{cal}}}{N_{\text{cal}}}(1{-}\alpha)-\frac{\eta}{\sqrt{N_{\text{cal}}}},$ (21)
where $N_{\text{cal}}=|\mathcal{D}_{\text{cal}}|$ and $\eta=\sqrt{\nicefrac{{\log(8)}}{{2}}}+\nicefrac{{1}}{{3}}$.
Proof.
The result follows directly from Theorem 1 from yang2024selection. ∎
3.3 Relation to other methods
@@ 4 Optimal Minimal Length Threshold
A key limitation of CPUL, like most well-known CP methods, is its tendency to produce overly narrow prediction intervals, particularly when the initial bounds $\hat{B}^{u}(x)$ and $\hat{B}^{l}(x)$ are tight. The length of the prediction intervals in Eq. (19) scales with the magnitude of $\hat{B}^{u}(x)-\hat{B}^{l}(x)$. Consequently, when the bounds are extremely tight, the constant adjustment of $t$ in the NCP framework becomes disproportionately large, leading empty prediction intervals and, in turn, under-coverage in such regions. Figure 2 illustrates this issue in the CPUL-lu setting. After calibrating the initial bounds the area around $x_{0}$ remains uncovered due to the empty intersection of the calibrated intervals. This behavior results in inefficient prediction intervals, as the method should over-cover the less confident regions (i.e., areas where $\hat{B}^{u}(x)-\hat{B}^{l}(x)$ is large). Hence, the regions with tightest initial bounds become most vulnerable and become under-covered.
Figure 2: Illustration of Paradoxical Miscoverage: the use of a constant offset (±$\pm t$) in the NCP construction results in an empty prediction interval in regions where $\hat{B}^{u}(x)-\hat{B}^{l}(x)$ is small.
sesia2020comparison proposed a variant, CQR-r, with
$\hat{C}_{t}^{\text{CQR-r}}(x)=\left[\hat{L}_{t}^{\text{CQR-r}}(x),\hat{U}_{t}^{\text{CQR-r}}(x)\right],\ \text{for }x\in\mathcal{X},$
where
$\hat{L}_{t}^{\text{CQR-r}}(x)=\hat{q}_{\alpha/2}(x)-t\left(\hat{q}_{1-\alpha/2}(x)-\hat{q}_{\alpha/2}(x)\right),$
@@
where $\Delta(x)\,{=}\,\hat{B}^{u}(x)\,{-}\,\hat{B}^{l}(x)$. It is easy to verify that $\{\bar{C}_{\ell,t}\}_{t\in\mathcal{T}}$ is a nested family, therefore satisfying the coverage guarantee given in (12).
The minimum length threshold $\ell$ may be specified by the user as a hyper-parameter, obtained as the solution of the optimization problem
@@
OMLT identifies regions where the given $\hat{B}^{u}$ and $\hat{B}^{l}$ are tightest, representing the most critical areas in CPUL, where even minor shrinking during calibration poses a high risk of significant undercoverage below the desired level of $1-\alpha$. OMLT retains the standard fixed-length adjustment of $t$ while introducing a threshold for the minimal allowed prediction interval length, effectively marking the boundary of high-risk regions. OMLT uses a hyperparameter $\ell$ that represents the minimal length of a prediction interval to enforce this threshold. For any given nested interval $\hat{C}_{t}(x)$, $x\in\mathcal{X}$ define
$t^{\prime}_{\ell}(x)=\inf_{t\in\mathcal{T}}\{t:\left|\hat{C}_{t}(x)\right|\geq\ell\}.$
For any $\hat{C}_{t}(x)$, define its corresponding nested interval leveraging the minimal-length threshold $\ell$ as
$\hat{C}_{t,\ell}^{\text{opt}}(x)=$ (28)
$\begin{cases}\hat{C}_{t}(x),&\hat{B}^{u}(x)-\hat{B}^{l}(x)>\ell,t\geq t^{\prime}_{\ell}(x);\\ \hat{C}_{t^{\prime}_{\ell}(x)}(x),&\hat{B}^{u}(x)-\hat{B}^{l}(x)>\ell,t<t^{\prime}_{\ell}(x);\\ [\hat{B}^{l}(x),\hat{B}^{u}(x)],&\hat{B}^{u}(x)-\hat{B}^{l}(x)\leq\ell.\end{cases}$
The minimum allowable range $\ell$, which defines the high-risk zone during calibration, is given by the optimization
$\min_{\ell\geq 0,t\in\mathcal{T}}$ $\mathbb{E}_{{X}}[\left|\hat{C}^{\text{opt}}_{t,\ell}({X})\right|]$ (29a)
[…]
The key idea underlying OMLT is that the size of the prediction interval should not be smaller than $\ell$, which prevents under-coverage when a prediction interval is too small. The only exception is for the tight regions, i.e., when $\hat{B}^{u}(x)-\hat{B}^{l}(x)\leq\ell$, in which case there is no need to enlarge $\hat{C}_{t,\ell}^{\text{opt}}(x)$ beyond $\ell$. The original upper and lower bounds can be confidently relied upon, as $[\hat{B}^{l}(x),\hat{B}^{u}(x)]$ already provides high-quality coverage.
Table 1: Summary of CP Interval Constructions
Method References Lower Bound Upper Bound
Split CP w/ $\hat{B}^{u}$ vovk2005algorithmic $\max\{\hat{B}^{u}+\tilde{Q}_{\frac{\alpha}{2}}^{u},\hat{B}^{l}\}$ $\hat{B}^{u}+\tilde{Q}^{u}_{1-\frac{\alpha}{2}}$
Split CP w/ $\hat{B}^{l}$ vovk2005algorithmic $\hat{B}^{l}+\tilde{Q}^{l}_{\frac{\alpha}{2}}$ $\min\{\hat{B}^{l}+\tilde{Q}^{l}_{1-\frac{\alpha}{2}},\hat{B}^{u}\}$
CQR romano2019conformalized $\hat{B}^{l}+{t}$ $\hat{B}^{u}-{t}$
CQR-r sesia2020comparison $\hat{B}^{l}+{t}(\hat{B}^{u}-\hat{B}^{l})$ $\hat{B}^{u}-{t}(\hat{B}^{u}-\hat{B}^{l})$
SFD CP sesia2023conformal $\max\big{\{}\hat{B}^{u}+\hat{Q}^{u}_{\frac{\alpha}{2}}-t,\hat{B}^{l}\big{\}}$ $\min\big{\{}\hat{B}^{l}+\hat{Q}^{u}_{1-\frac{\alpha}{2}}+t,\hat{B}^{u}\big{\}}$
CPUL Algorithm 1 $\hat{L}^{**}_{{t}}$ $\hat{U}^{**}_{{t}}$
CPUL-OMLT Problem (29) + CPUL $\hat{L}^{\text{opt},**}_{{t},\ell}$ $\hat{U}^{\text{opt},**}_{{t},\ell}$
  • $t$ is updated after calibration using (12) for all methods except Split CP, where the calibration results are provided by $\tilde{Q}^{l}_{\alpha}$, $\tilde{Q}^{l}_{1-\alpha}$, $\tilde{Q}^{u}_{\alpha}$, and $\tilde{Q}^{u}_{1-\alpha}$.
@@ 5 Experiments
The performance of CPUL and CPUL-OMLT are evaluated to demonstrate their ability to achieve valid coverage while producing narrower prediction intervals. The experiments focus on UQ for the optimal value of economic dispatch problems, detailed in Appendix A.1. UQ is conducted given lower and upper bounds derived from primal-dual optimization proxy models chen2023end; qiu2024dual; klamkin2024dual; chen2024real. Namely, the dual proxy provides valid lower bounds ($\hat{B}^{l}$), and the primal proxy provides valid upper bounds $\hat{B}^{u}$).
Further details on these optimization proxies are provided in Appendices A.1 and A.3.
Table 2: Performance Comparison of CP Methods
$\alpha=5\%$ $\alpha=10\%$
UQ Method PICP (%) Size (%) PICP (%) Size (%)
89_pegase
$[\hat{B}^{l},\hat{B^{u}}]$ 100.0 (0.00) 0.410 (0.016) 100.0 (0.00) 0.410 (0.016)
Split CP w/ $\hat{B}^{l}$ 95.02 (0.40) 0.242 (0.026) 90.15 (0.60) 0.197 (0.007)
Split CP w/ $\hat{B}^{u}$ 95.02 (0.43) 0.230 (0.006) 90.02 (0.54) 0.205 (0.007)
SFD CP 95.97 (0.27) 0.215 (0.009) 91.23 (0.56) 0.188 (0.007)
CQR 95.95 (1.05) 0.398 (0.018) 91.40 (1.80) 0.389 (0.018)
CQR-r 95.37 (0.46) 0.359 (0.020) 90.71 (0.72) 0.332 (0.020)
CPUL 95.97 (0.27) 0.215 (0.009) 91.23 (0.56) 0.188 (0.007)
CPUL-OMLT 95.97 (0.27) 0.215 (0.009) 90.28 (0.51) 0.187 (0.008)
118_ieee
$[\hat{B}^{l},\hat{B^{u}}]$ 100.0 (0.00) 0.281 (0.085) 100.0 (0.00) 0.281 (0.085)
Split CP w/ $\hat{B}^{l}$ 95.30 (0.48) 0.230 (0.148) 90.34 (0.63) 0.206 (0.146)
Split CP w/ $\hat{B}^{u}$ 95.39 (0.26) 0.127 (0.005) 90.00 (0.50) 0.105 (0.004)
SFD CP 95.19 (0.30) 0.162 (0.008) 90.10 (0.50) 0.135 (0.013)
CQR 95.57 (0.69) 0.205 (0.147) 90.08 (2.75) 0.190 (0.147)
CQR-r 95.62 (0.78) 0.196 (0.168) 90.88 (1.33) 0.180 (0.167)
CPUL 95.45 (0.30) 0.127 (0.005) 90.02 (0.46) 0.105 (0.004)
CPUL-OMLT 95.41 (0.23) 0.126 (0.006) 90.01 (0.48) 0.103 (0.007)
1354_pegase
$[\hat{B}^{l},\hat{B^{u}}]$ 100.0 (0.00) 3.949 (3.807) 100.0 (0.00) 3.949 (3.807)
Split CP w/ $\hat{B}^{l}$ 94.64 (0.32) 1.324 (1.212) 89.60 (0.54) 1.308 (1.160)
Split CP w/ $\hat{B}^{u}$ 94.66 (0.37) 2.680 (2.724) 89.40 (0.49) 1.570 (0.853)
SFD CP 95.01 (0.25) 2.324 (2.037) 90.22 (0.27) 1.456 (0.609)
CQR 94.75 (0.36) 3.552 (3.943) 90.13 (0.43) 3.501 (3.939)
CQR-r 95.37 (0.65) 3.406 (3.947) 91.57 (1.32) 3.271 (3.997)
CPUL 94.62 (0.33) 1.321 (1.215) 89.65 (0.56) 1.306 (1.162)
CPUL-OMLT 94.66 (0.42) 1.156 (0.896) 89.69 (0.49) 1.037 (0.890)
@@
Experiments are performed over three datasets: 118_ieee ieee_pstca, 1354_pegase pegase, and 89_pegase pegase, corresponding to power grids of different sizes.
Random shuffling and splitting of data samples is performed 10 times separately for each dataset. Detailed descriptions of these datasets are provided in Section A.2. The evaluation metrics are presented as the mean and standard deviation across the 10 experiments.
@@
The performance of CPUL and CPUL-OMLT is compared against several CP baselines; Table 1 presents an overview of each method considered in the experiments and the form of the corresponding prediction intervals (third and fourth columns). Recall that, unless specified otherwise, all prediction intervals are strengthened as per Eq. (15). The CP baselines include: Split CP (Eq. 7), using either $\hat{B}^{l}$ or $\hat{B}^{u}$ as base predictors; CQR and its variant CQR-r, wherein $\hat{B}^{l}$ and $\hat{B}^{u}$ are treated as the initial lower and upper quantile regressors in the CQR construction; the SFP CP approach of sesia2023conformal. Note that the latter matches the $\hat{C}^{\text{ul}}$ construction. Additional implementation details are provided in Appendix A.4.1.
@@
All methods are evaluated on a held out test set $\mathcal{D}_{\text{test}}\,{=}\,(X_{i},Y_{i})_{i\in\mathcal{I}_{\text{test}}}$, using two evaluation metrics: interval length and coverage. At a given confidence level $1\,{-}\,\alpha$, shorter interval lengths are considered more desirable. This is captured by the expected size of prediction interval $\hat{C}$, defined as
$\hat{E}_{X}(\hat{C})=\frac{1}{|\mathcal{I}_{\text{test}}|}\sum_{i\in\mathcal{I}_{\text{test}}}\frac{\left|\hat{C}(X_{i})\right|}{\left|Y_{i}\right|},$
where $\hat{C}(X_{i})$ denotes the prediction interval for $X_{i}$. The paper considers the scaled interval length $|\hat{C}(X_{i})|/|Y_{i}|$, expressed as a percentage, rather than absolute interval length $|\hat{C}(X_{i})|$, to account for the possibly large range of values taken by $Y$.
The evaluation metrics also include the Prediction Interval Coverage Percentage (PICP) that represents the proportion of true values that fall within the predicted intervals across the test dataset. Mathematically, the PICP is computed as:
$\text{PICP}=\frac{1}{|\mathcal{I}_{\text{test}}|}\sum_{i\in\mathcal{I}_{\text{test}}}\mathbf{1}_{\hat{C}(X_{i})}(Y_{i}).$
Finally, recall that average and standard deviation across 10 runs are reported for every metric.
@@
Table 2 and Figure 3 present the numerical performance of the various methods. Table 2 presents, for every dataset, the average and standard deviation of each method’s interval length and coverage, both expressed as a percentage. The table reports results for $\alpha=5\%$ and $\alpha=10\%$; additional $\alpha$ values are reported in Table 4 in Appendix B. Figure 3 displays the interaction between each method’s interval length and coverage, across a broader range of $\alpha$ values. This provides a more global view of each method’s overall performance.
@@
The results in Table 2 demonstrate that all CP methods achieve satisfactory coverage levels, i.e., test coverage is typically close to $1-\alpha$, which is expected given that all methods are properly calibrated. Moreover, all prediction intervals yield a substantial reduction in size compared to the original prediction $[\hat{B}^{l},\hat{B}^{u}]$. For instance, a reduction of only 5% in coverage (i.e., 95% prediction intervals) can yield up to a two- and three-fold reduction in interval size on the 89_pegase and 1354_pegase, respectively. This demonstrates the value of using UQ techniques to provide more actionable information to practitioners.

### v2

@@ 3.1 Exploiting valid bounds in NCP
Recall that $\hat{B}^{l},\hat{B}^{u}$ provide valid lower and upper bounds on the target variable, i.e., $\hat{B}^{l}(X)\,{\leq}\,Y\,{\leq}\,\hat{B}^{u}(X)$ always holds. This suggests a simple procedure for strengthening $\hat{C}$ without loss of coverage (as shown by Proposition 3.1) into
@@
$\hat{C}(x)\,{\cap}\,\big{[}\hat{B}^{l}(x),\hat{B}^{u}(x)\big{]},\forall x\in\mathcal{X}$. Then, $\mathbb{P}\big{(}Y_{N+1}\,{\in}\,\tilde{C}(X_{N+1})\big{)}=\mathbb{P}\big{(}Y_{N+1}\,{\in}\,\hat{C}(X_{N+1})\big{)}=1{-}\alpha.$appendixproof} It suffices to prove that $\mathbb{P}\left(Y\in\tilde{C}(X)\right)=\mathbb{P}\left(Y\in\hat{C}(X)\right)$. First note that
@@
which concludes the proof. It is important to note that Proposition 3.1 holds irrespective of how the original prediction interval $\hat{C}$ is obtained. Theorem 3.1 shows that, if the calibration step is performed using the NCP framework, then there is no loss of performance whether the strengthening is performed before or after calibration. {theoremrep} Consider a family of nested prediction sets $\{\hat{C}_{t}\}_{t\in\mathcal{T}}$, where $\mathcal{T}\,{\subseteq}\,\mathbb{R}$, and let $\hat{\tau}$ be obtained following the NCP calibration step as per Eq. (11). Next, define the family of nested strengthened intervals $\{\tilde{C}_{t}\}_{t\in\mathcal{T}}$, where $\forall x\in\mathbb{R}^{d},\forall t\in\mathcal{T},\tilde{C}_{t}(x)=\hat{C}_{t}(x)\cap\big{[}\hat{B}^{l}(x),\hat{B}^{u}(x)\big{]},$ and let $\tilde{\tau}$ be obtained from Eq. (11). Then, $\hat{\tau}=\tilde{\tau}$ and $\forall x\in\mathbb{R}^{d},$
$\hat{C}_{\hat{\tau}}(x)\cap[\hat{B}^{l}(x),\hat{B}^{u}(x)]$ $=\tilde{C}_{\tilde{\tau}}(x).$ (15)
@@
which then yields $\hat{\tau}=\tilde{\tau}$ by unicity of the infimum, and $\hat{C}_{\hat{\tau}}(X)\cap[\hat{B}^{l}(X),\hat{B}^{u}(X)]=\tilde{C}_{\tilde{\tau}}(X)$ follows immediately. Theorem 3.1 allows to decouple the calibration step from the strengthening. Thereby, one can perform the calibration step without access to $\hat{B}^{l}$ or $\hat{B}^{u}$, without any loss of coverage. This is particularly valuable when $\hat{B}^{l}$ or $\hat{B}^{u}$ are not available during training/calibration due to, e.g., privacy concerns.
@@ 3.2 The CPUL approach
Note that $L^{l}$, $L^{u}$, $U^{l}$, $U^{u}$ do not provide valid lower nor upper bounds on the target variable ($Y$). The motivation for adjusting the initial predictors $\hat{B}^{l},\hat{B}^{u}$ using quantiles $\hat{Q}^{l},\hat{Q}^{u}$ is to account for the heteroskedascitity of $\hat{r}^{l},\hat{r}^{u}$. Also note that the construction in Eq. (18) only requires evaluating the residuals $\hat{r}^{l},\hat{r}^{u}$ on the training set, and extracting their quantiles.
@@
This construction is illustrated in Figure 1(a). Each family is then conformalized using the NCP procedure on the calibration set, and the model with smallest width is selected.
Algorithm 1 summarizes the proposed CPUL method, which proceeds as follows. First, the dataset $\mathcal{D}$ is split into training and calibration sets (line 1). Then, lower and upper bound predictors (i.e., $\hat{B}^{l},\hat{B}^{u}$) are trained using the training set (line 2), after which quantiles of the corresponding residuals are evaluated on the training set (line 3). Next, four families of nested intervals are constructed following Eq. (19), each of which is calibrated using the NCP framework (Eq. (11)) and the calibration set $\mathcal{D}_{\text{cal}}$ (line 4). The final step of the algorithm (line 5) selects, among the four models $\hat{C}^{\text{ll}},\hat{C}^{\text{lu}},\hat{C}^{\text{ul}},\hat{C}^{\text{uu}}$, the one with smallest average width over the calibration set (as an estimator of $\mathbb{E}\big{|}\hat{C}(X_{N+1})\big{|}$). It is important to note that the model selection step is not performed on a per-sample basis. Rather, Algorithm 1 selects one variant among $\hat{C}^{\text{ll}},\hat{C}^{\text{lu}},\hat{C}^{\text{ul}},\hat{C}^{\text{uu}}$, which is then used across the entire test set. For instance, if $\hat{C}^{\text{lu}}_{\tau^{\text{lu}}}$ is selected, then $\hat{C}^{**}\,{=}\,\hat{C}^{\text{lu}}_{\tau^{\text{lu}}}$ and the CPUL prediction interval is $\hat{C}^{**}(X_{N+1})\,{=}\,\hat{C}^{\text{lu}}_{\tau^{\text{lu}}}(X_{N+1})$. Theorem 1 provides theoretical guarantees on the coverage of $\hat{C}^{**}$.
Theorem 1.
Assume that $\{(X_{1},Y_{1}),...,(X_{N+1},Y_{N+1})\}$ are i.i.d. samples, and let $\hat{C}^{**}$ be the CPUL model selected by Algorithm 1. Define $N_{\text{cal}}=|\mathcal{D}_{\text{cal}}|$ and $\eta=\sqrt{\nicefrac{{\log(8)}}{{2}}}+\nicefrac{{1}}{{3}}$. Then
$\mathbb{P}\left(Y_{N+1}\in\hat{C}^{**}(X_{N+1})\right)\geq\frac{1{+}N_{\text{cal}}}{N_{\text{cal}}}(1{-}\alpha)-\frac{\eta}{\sqrt{N_{\text{cal}}}}.$ (20)
Proof.
The result follows directly from Theorem 1 from yang2024selection. ∎
3.3 Relation to other methods
@@ 4 Optimal Minimal Length Threshold
A key limitation of CPUL, like most well-known CP methods, is its tendency to produce overly narrow prediction intervals, particularly when the initial bounds $\hat{B}^{u}(x)$ and $\hat{B}^{l}(x)$ are tight. The length of the prediction intervals in Eq. (18) scales with the magnitude of $\hat{B}^{u}(x)-\hat{B}^{l}(x)$. Consequently, when the bounds are extremely tight, the constant adjustment of $t$ in the NCP framework becomes disproportionately large, leading empty prediction intervals and, in turn, under-coverage in such regions. Figure 1(b) illustrates this issue in the CPUL-lu setting. After calibrating the initial bounds the area around $x_{0}$ remains uncovered due to the empty intersection of the calibrated intervals. This behavior results in inefficient prediction intervals, as the method should over-cover the less confident regions (i.e., areas where $\hat{B}^{u}(x)-\hat{B}^{l}(x)$ is large). Hence, the regions with tightest initial bounds become most vulnerable and become under-covered. sesia2020comparison proposed a variant, CQR-r, with $\hat{C}_{t}^{\text{CQR-r}}(x)=\left[\hat{L}_{t}^{\text{CQR-r}}(x),\hat{U}_{t}^{\text{CQR-r}}(x)\right],\ \text{for }x\in\mathcal{X},$ where
$\hat{L}_{t}^{\text{CQR-r}}(x)=$ $\hat{q}_{\alpha/2}(x)-t\left(\hat{q}_{1-\alpha/2}(x)-\hat{q}_{\alpha/2}(x)\right);$
@@
where $\Delta(x)\,{=}\,\hat{B}^{u}(x)\,{-}\,\hat{B}^{l}(x)$. It is easy to verify that $\{\bar{C}_{\ell,t}\}_{t\in\mathcal{T}}$ is a nested family, therefore satisfying the coverage guarantee given in (11). The minimum length threshold $\ell$ may be specified by the user as a hyper-parameter, obtained as the solution of the optimization problem
@@ 5 Experiments
The performance of CPUL and CPUL-OMLT are evaluated to demonstrate their ability to achieve valid coverage while producing narrower prediction intervals. The experiments focus on UQ for the optimal value of economic dispatch problems, detailed in Appendix A.1. UQ is conducted given lower and upper bounds derived from primal-dual optimization proxy models chen2023end; qiu2024dual; klamkin2024dual; chen2024real. Namely, the dual proxy provides valid lower bounds ($\hat{B}^{l}$), and the primal proxy provides valid upper bounds $\hat{B}^{u}$). Further details on these optimization proxies are provided in Appendices A.1 and A.3.
Table 1: Performance Comparison of CP Methods ($\alpha=10\%$)
89_pegase 118_ieee 1354_pegase
UQ Method PICP (%) Size (%) PICP (%) Size (%) PICP (%) Size (%)
$[\hat{B}^{l},\hat{B^{u}}]$ 100.0 (0.00) 0.410 (0.016) 100.0 (0.00) 0.281 (0.085) 100.0 (0.00) 3.949 (3.807)
Split CP w/ $\hat{B}^{l}$ 90.15 (0.60) 0.197 (0.007) 90.34 (0.63) 0.206 (0.146) 89.60 (0.54) 1.308 (1.160)
Split CP w/ $\hat{B}^{u}$ 90.02 (0.54) 0.205 (0.007) 90.00 (0.50) 0.105 (0.004) 89.40 (0.49) 1.570 (0.853)
SFD CP 91.23 (0.56) 0.188 (0.007) 90.10 (0.50) 0.135 (0.013) 90.22 (0.27) 1.456 (0.609)
CQR 91.40 (1.80) 0.389 (0.018) 90.08 (2.75) 0.190 (0.147) 90.13 (0.43) 3.501 (3.939)
CQR-r 90.71 (0.72) 0.332 (0.020) 90.88 (1.33) 0.180 (0.167) 91.57 (1.32) 3.271 (3.997)
CPUL (ours) 91.23 (0.56) 0.188 (0.007) 90.02 (0.46) 0.105 (0.004) 89.65 (0.56) 1.306 (1.162)
CPUL-OMLT (ours) 90.28 (0.51) 0.187 (0.008) 90.01 (0.48) 0.103 (0.007) 89.69 (0.49) 1.037 (0.890)
    For each dataset, the three shortest intervals are colored blue, while the three largest intervals are colored red.
@@
Experiments are performed over three datasets: 118_ieee ieee_pstca, 1354_pegase pegase, and 89_pegase pegase, corresponding to power grids of different sizes. For each dataset, samples are randomly shuffled and split 10 times. Dataset details are provided in Section A.2. Evaluation metrics reported are the mean and standard deviation across these 10 runs.
@@
The performance of CPUL-OMLT is compared against several CP baselines. Recall that, unless specified otherwise, all prediction intervals are strengthened as per Eq. (14). The CP baselines include: Split CP (Eq. 6), using either $\hat{B}^{l}$ or $\hat{B}^{u}$ as base predictors; CQR and its variant CQR-r, wherein $\hat{B}^{l}$ and $\hat{B}^{u}$ are treated as the initial lower and upper quantile regressors in the CQR construction; an adapted SFP CP approach of sesia2023conformal. Note that the latter matches the $\hat{C}^{\text{ul}}$ construction. Additional implementation details are provided in Appendix A.4.1.
@@
All methods are evaluated on a held out test set $\mathcal{D}_{\text{test}}\,{=}\,(X_{i},Y_{i})_{i\in\mathcal{I}_{\text{test}}}$, using two evaluation criteria: coverage and interval length. The former is evaluated via the Prediction Interval Coverage Percentage (PICP), which measures the proportion of true values contained within the predicted intervals on the test set, i.e., $\text{PICP}=\frac{1}{|\mathcal{I}_{\text{test}}|}\sum_{i\in\mathcal{I}_{\text{test}}}\mathbf{1}_{\hat{C}(X_{i})}(Y_{i}).$
At a given confidence level $1\,{-}\,\alpha$, shorter interval lengths are considered more desirable. This is captured through expected normalized length of prediction interval $\hat{C}$, defined as $\hat{E}_{X}(\hat{C})=|\mathcal{I}_{\text{test}}|^{-1}\sum_{i\in\mathcal{I}_{\text{test}}}\big{(}\left|Y_{i}\right|^{-1}|\hat{C}(X_{i})|\big{)}$, where $Y_{i}$ and $\hat{C}(X_{i})$ denote the (true) optimal value and the prediction interval for sample $i$. The paper considers the scaled interval length $|\hat{C}(X_{i})|/|Y_{i}|$, expressed as a percentage, rather than absolute interval length $|\hat{C}(X_{i})|$, to account for the possibly large range of values taken by $Y$. This metric is routinely used in the optimization literature chen2023end; vivas2020systematic; wang2009comparison, where it is referred to as optimality gap. All metrics are reported as the mean and standard deviation over 10 runs.
Experiment Results
@@
Table 1 and Figure 2 present the numerical performance of the various methods. Table 1 presents, for every dataset, the average and standard deviation of each method’s interval length and coverage, both expressed as a percentage. The table reports results for $\alpha=10\%$; additional $\alpha$ values are reported in Table 3 in Appendix B. Figure 2 displays the interaction between each method’s interval length and coverage, across a broader range of $\alpha$ values. This provides a more global view of each method’s overall performance.
@@
The results in Table 1 demonstrate that all CP methods achieve satisfactory coverage levels, i.e., test coverage is typically close to $1-\alpha$, which is expected given that all methods are properly calibrated. Moreover, all prediction intervals yield a substantial reduction in size compared to the original prediction $[\hat{B}^{l},\hat{B}^{u}]$. For instance, a reduction of only 10% in coverage (i.e., 90% prediction intervals) can yield up to a two- and three-fold reduction in interval size. This demonstrates the value of using UQ techniques to provide more actionable information to practitioners.